# Top-down machine learning of coarse-grained protein force-fields

### v1

@@ 3.3 The Neural Network Potential is comparable to other methods in de-novo structure prediction
To benchmark our model (G-NNP) against other methods, we calculated the average root mean square deviation (RMSD) of the most probable macrostate derived from the simulations. The results are presented in Table 2. We compared our findings with two web servers employing coarse-grained methods for protein folding, UNRES 50 and CABS-fold 51, as well as the only other method utilizing differentiable molecular simulations to learn coarse-grained parameters, DMS 34. For the CABS-fold and DMS methods, the results were obtained from the paper by Greener et al. 34, as the CABS-fold server was not operational during our analysis. For the UNRES method, models were generated on their web server using the parameters provided in the MREMD structure prediction example from their tutorial.
@@
Chignolin 2.7 4.8 4.8 4.8
Trp-cage 5.6 2.7 3.5 5.7
BBA 3.6 7.2 7.9 2.5
Villin 7.4 6.4 11.5 8.2
BBL 11.6 9.1
WW-domain 8.7 8.1
Protein B 7.1 10.9
Protein G 11.7 9.6
NTL9 8.3 8.3
Homeodomain 5.3 8.3
a3d 8.9 14.0
$\lambda$-repressor 9.2 11.5
@@
Another aspect of our method is its capacity to illustrate not only the end conformation, as current protein structure prediction methods48, 52, 53, but also the pathway the protein traverses towards it. This aspect could provide a more comprehensive understanding of protein dynamics, and in combination with additional reference data, it could eventually predict both structure and folding pathways.

### v4

@@ 3.3 The Neural Network Potential is comparable to other methods in de-novo structure prediction
While AlphaFold258 has indeed revolutionized the domain with its ability to predict native state models, it largely depends on multiple sequence alignments (MSAs) during the initial stages of native state prediction. In contrast, our approach does not require the utilization of such information, positioning it as potentially more versatile, particularly in contexts where obtaining MSAs is a challenge. Additionally, our model holds the potential of not only predict the native structure but also the whole folding process.
@@
To benchmark our model (G-NNP) against other methods, we calculated the average root mean square deviation (RMSD) of the most probable macrostate derived from the simulations. The results are presented in Table 2. We compared our findings with two web servers employing coarse-grained methods for protein folding, UNRES 60 and CABS-fold 61, as well as the only other method utilizing differentiable molecular simulations to learn coarse-grained parameters, DMS 44. For the CABS-fold method, as the server was not operational during our analysis, the results for Chignolin, Trp-cage, BBA, and Villin were obtained from the paper by Greener et al. 44, for DMS we run the predictions using the same settings they use in their paper and for the UNRES method, models were generated on their web server using the parameters provided in the MREMD structure prediction example from their tutorial.
@@
Chignolin 2.7 4.8 4.8 5.28
Trp-cage 5.6 2.7 3.5 5.47
BBA 3.6 7.2 7.9 7.54
Villin 7.4 6.4 11.5 7.21
BBL 8.68 11.6 6.26
WW-domain 9.07 8.7 8.46
Protein B 8.88 7.1 6.56
Protein G 11.59 11.7 9.76
NTL9 9.58 8.3 7.34
Homeodomain 6.82 5.3 6.17
a3d 11.81 8.9 3.73
$\lambda$-repressor 12.48 9.2 11.2
@@
Another aspect of our method, and the ones we have used as a benchmarks (UNRES, CABS-fold and DMS) is their capacity to illustrate not only the end conformation, as current protein structure prediction methods 58, 62, 63, but also the pathway the protein traverses towards it. This aspect could provide a more comprehensive understanding of protein dynamics, and in combination with additional reference data, it could eventually predict both structure and folding pathways.
It is worth noting that our current model does not fully encapsulate the comprehensive reproduction of the entire conformational landscape at this stage. Despite this, we envision our method as a significant stepping stone paving the way for future advancements in the field. Looking ahead, we perceive the potential of this approach to be used as a pre-training stage that can be trained on large amounts of proteins to capture relevant information. Subsequently, it could be combined with active learning strategies to learn the exact forces in sampled conformations, thus more accurately mirroring the thermodynamics, or used as a foundation model that can be fine-tuned for specific downstream tasks.
@@ 4 Conclusions
Associated Content
Supporting Information available. The SI provides detailed mathematical formulations for the prior energy terms used in the model, as well as the architecture and equations for the Graph Neural Networks. It also includes training curves for different potentials (Figures S1-S2), a comparative analysis of the free energy landscape across initial TICA dimensions (Figure S3), and a table listing the sequences of fast-folding proteins studied (Table S1). Additional tables outline the hyperparameters used for neural network training (Table S2) and a comparison of MD simulation speeds for different proteins using different NNPs (Table S3).